OpenStudy (anonymous):

int[1, 3] 1/(x^3+x) dx

OpenStudy (anonymous):

\[\int\limits_{1}^{3} \frac{1}{(x^{3}+x)} dx\]

OpenStudy (amistre64):

partial decomp by chance?

OpenStudy (amistre64):

\[\frac{1}{x(x^2+1)}\]

OpenStudy (anonymous):

\[\frac{ 1 }{ x } - \frac{ x }{ x^{2}+1 }\]

OpenStudy (anonymous):

how would i integrate \[\frac{ x }{ x^{2}+1 }\] ?

OpenStudy (anonymous):

\[\int\limits_{}(1/x) dx=\log(x)\]

OpenStudy (amistre64):

Dx[x^2+1] = 2x, so your missing a 2. multiply by 2/2 and borrow it from that

OpenStudy (anonymous):

\[\int\limits_{}^{}x/(x ^{2}+1) dx\] put x^2=t 2x dx= dt the integral reduces to

OpenStudy (anonymous):

\[(1/2)\int\limits_{}^{}1/(1+t) dt\]

OpenStudy (anonymous):

any confusion?/

OpenStudy (anonymous):

yeah partial decomp or long division

OpenStudy (anonymous):

ahh, thank you! i should be fine from here on!

OpenStudy (anonymous):

you are welcome....!!!!